Hi
Tom traveled for 2 hours at 10mph then stopped for 1 hour then continued to travel at the same speed.Gerald started his travel along the same route 3.75 hours laterthan tom at 32mph.
how far from the starting point will gerald be when he catches up with tom.
thanks

Found 2 solutions by richwmiller, greenestamps:
Answer by richwmiller(17219)

Tom traveled for 2 hours at 10 mph then stopped for 1 hour then continued to travel at the same speed. Gerald started his travel along the same route 3.75 hours later than Tom at 32 mph.
How far from the starting point will Gerald be when he catches up with Tom?
In the first 2 hours, Tom went 20 miles
The next hour he rested
3.75-3=.75 hr.
Then he went .75 *10=7.50 miles
Tom had a headstart of 27.5 mi.
10x+27.5=32x
27.5=22x
x=1.25 hrs
32*1.25=40
They met at the 40 mi marker
check
1.25*10=12.5 mi
27.5+12.5.=40

Answer by greenestamps(13199)



Since Gerald started 3.75 hours after Tom, and since Tom rested during one of those hours, the distance between them when Gerald starts is 2.75(10) = 27.5 miles.

The difference in their speeds is 32-10 = 22mph.

The time required for Gerald to catch up with Tom is the difference in the distances, divided by the difference in the speeds: 27.5/22 = 1.25 hours.

Gerald traveling as 32mph for 1.25 hours will travel 32(1.25) = 40 miles.

ANSWER: Gerald will catch up with Tom 40 miles from the starting point.

CHECK: When Gerald has traveled for 1.25 hours, Tom has traveled 2.75+1.25 = 4 hours; 4 hours at 10mph equals 40 miles.